OpenStudy (anonymous):

Jackson has 5 more CDs than Amal. They have a total of 95 CDs. How many CDs does Amal have?

OpenStudy (mysesshou):

J = A + 5 (eq1) J + A = 95 (eq2) Solve the 2 equations, 2 variables. You can plug in (eq1) into (eq2) (A + 5) + A = 95 subtract 5 from each side gives 2*A = 90 divide both sides by 2 A = 45 Therefore, Amel has 45 and Jackson has 50. Hope this helps. :)

OpenStudy (anonymous):

it helps thanks!!!